OpenStudy (anonymous):

To find out how long it will take for $500 to double when invested at %6 annual interest compounded twice a year, you solve the following equation: 1000=500(1+.06/2)^2t. Use complete sentences to describe each step of math needed to solve this equation.

OpenStudy (amistre64):

solve for "t"

OpenStudy (anonymous):

how would you do that?

OpenStudy (amistre64):

A = P(1+b)^nt A/P = (1+b)^nt log(A/P) = nt log(1+b) log(A/P) ------------ = t n log(1+b)

OpenStudy (anonymous):

How would I plug the question into that

OpenStudy (amistre64):

by comparing similar parts and substituting in the values into the placeholders ..... like normal

OpenStudy (anonymous):

If you don't mind, can you show me?

OpenStudy (amistre64):

I did show you .... im not sure what else that i can do to make it any simpler compare the first equation i used with the equation given and youll be able to define A, P, b and n with them to replace with

OpenStudy (amistre64):

this is your given specific equation 1000=500(1+.06/2)^2t I made the values general to keep the workings simple to follow A = P(1 + b)^nt compare these to define the values for a specific result 1000 = 500 (1+.06/2) ^ 2t A = P (1 + b ) ^ nt

OpenStudy (anonymous):

does 2.19=t

OpenStudy (amistre64):

id have to check it on a calculator ..... ln(2)/(2ln(1.03)) = 11.something